********************************************************************************* C*-NEWS 42 1*06*1994/31*12*1994 *********************************************************************************

C*-94/P.3490 CHU C.-H., KUSUDA M. On Factor States of C*-Algebras and their Extensions Goldsmiths College, Univ. London, London SE14 6NW (U.K.)

C*-94/P.3491 KOSAKI H. Sector Theory and Automorphisms for Factor-Subfactor Pairs Graduate School of Math., Kyushu Univ., Fukuoka 810 (Japan)

C*-94/P.3492 JORGENSEN J.E.T., PEDERSEN S. Local Harmonic Analysis for Domains in R^{n} of Finite Measure Dpt. Math., Univ. Iowa. Iowa City IA 52242 (USA) E-MAIL : jorgen@math.uiowa.edu

spedersen@desire.wright.edu

C*-94/P.3493 OCNEANU A. Chirality for Operator Algebras Dpt. Math., Pennsylvania State Univ., University Park PA 16802 (USA)

C*-94/P.3494 FRANK M., MANUILOV V.M. Diagonalizing ”Compact” Operators on Hilbert W*-Modules Univ. Leipzig, Fak. Math. Inform., Math. Inst., Augustusplatz 10 04109 Leipzig (Allemagne)

C*-94/P.3495 BEZUGLYI S.I., GOLODETS V.YA. Dynamical Entropy for Z^{n}-Actions and Bogoliubov Automorphisms of the CAR-Algebra Erwin Schrodinger Int. Inst. for Math. Physics Inst. Low Temperature Phys. Engin., 47 Lenin Ave., 310164 Kharkov (Ukraine)

C*-94/P.3496 GUIDO D., LONGO R. An Algebraic Spin and Statistics Theorem. I Dipart. di Matem., Univ. di Roma Tor Vergata, Vialle della Ricerca Scientifica, 00133 Roma (Italy) E-MAIL : guido@mat.utovrm.it

longo@mat.utovrm.it

C*-94/P.3499 SCHMUDGEN K. Classification of Bicovariant Differential Calculi on Quantum General Linear Groups Univ. Leipzig, Fak. Math. Inform., Math. Inst., Augustusplatz 10 04109 Leipzig (Allemagne) E-MAIL : schmuedgen@mathematik.uni-leipzig.d400.de

C*-94/P.3498 EVANS D.E., KAWAHIGASHI Y. On Ocneanu’s Theory of Asymptotic Inclusions for Subfactors, Topological Quantum Field Theories and Quantum Doubles (Revised) Dpt. Math., Univ. Wales Swansea, Singleton Park, Swansea SA2 8PP, Wales (U.K.) E-MAIL : yasuyuki@tansei.cc.u-tokyo.ac.jp

C*-94/P.3499 GOLODETS V.Y., SOKHET A.M. Cocycles of Type III Transformation Group and AT Property for the Double Mackey Action Erwin Schrodinger Intern. Inst. Math. Sciences, Pasteurgasse 6/7 A-1090 Wien (Austria)

C*-94/P.3500 CONNES A., ROVELLI C. Von Neumann Algebra Automorphisms and Time-Thermodynamics Relation in General Covariant Quantum Theories IHES, 35, Route de Chartres, 91440 Bures sur Yvette (France)

C*-94/P.3501 WINSLOW C. The Flow of Weights in Subfactor Theory Math. Inst., Kobenhavns Universitet, 2100 Copenhagen 0 (Denmark)

C*-94/P.3502 CURTO R.E., FIALKOW L.A. Solution of the Truncated Complex Moment Problem for Flat Data Dpt. Math., Univ. Iowa. Iowa City IA 52242 (USA) E-MAIL : jorgen@math.uiowa.edu

fialkow@snyewvm.bitnet C*-94/P.3503 ELLIOTT G.A., RORDAM M. Classification of Certain Infinite Simple C*-Algebras. II Inst. Mat. of Dat., Odense Univ., Campusvej 55, DK-5230 Odense M (Denmark) E-MAIL : mikael@imada.ou.dk

C*-94/P.3504 HAAGERUP U., RORDAM M. Perturbations of the Rotation C*-Algebras and of the Heisenberg Commutation Relation Inst. Mat. of Dat., Odense Univ., Campusvej 55, DK-5230 Odense M (Denmark) E-MAIL : haagerup@imada.ou.dk

mikael@imada.ou.dk

C*-94/P.3505 AYUPOV SH.A., FERLEGER S.V.,

SUKOCHEV F.A.

Are Non-Associative L^{P} -Spaces Really Non-Associative ? IRMA, Univ. Louis Pasteur, 7 rue R. Descartes, 67084 Strasbourg (France) E-MAIL : irma@math.u-strasbg.fr

C*-94/P.3506 EXEL R. Twisted Partial Actions. A Classification of Stable C*-Algebraic Bundles (Preliminary Version) Depart. de Matend., Univ. de Sao Paulo, C.P. 20570, 01452-990 Sao Paulo (Brasil) E-MAIL : exel@ime.usp.br

C*-94/P.3507 JORGENSEN P.E.T. Scattering Theory for Orthogonal Wavelets Dpt Math., Univ. of Iowa, Iowa City IA 52242 (USA) E-MAIL : jorgen@math.uiowa.edu

C*-94/P.3508 CURTO R.E., FIALKOW L.A. Solution of the Truncated Complex Moment Problem for Flat Data Dpt Math., Univ. of Iowa, Iowa City IA 52242 (USA)

C*-94/P.3509 LORING T.A., PEDERSEN G.K.

Corona Extendibility and Asymptotic Multiplicativity Univ. Copenhagen, Inst. Math., 5 Universitetsparken, DK-2100 Copenhagen 0 (Denmark) E-MAIL : spedersen@desire.wright.edu

C*-94/P.3510 LORING T.A. Perturbation Querstions in the Cuntz Picture of K-Theory Dpt. Math. Stat., Univ. New Mexico, Alberquerque, NM 87131 (USA)

C*-94/P.3511 KAWAHIGASHU Y. Paragroups as Quantized Galois Groups for Subfactors Dpt. Math. Sciences, Univ. Tokyo, Hongo, Tokyo 113 (Japan) E-MAIL : yasuyuki@tansei.cc.u-tokyo.ac.jp)

C*-94/P.3512 AYUPOV SH.A., YADGOROV N.J. Geometry of the State Spaces in Quantum Probability Inst. Math., Usk. Acad. Sci., F-Hodjaev str. 29, Tashkent 143 (700143 Rep. of Usbekistan)

C*-94/P.3513 DOPLICHER S., FREDENHAGEN K.,

ROBERT J.E.

The Quantum Structure of Spacetime at the Planck Scale and Quantum Fields Dip. di Matem., Univ. di Roma La Sapienza, I-00185 Roma (Italy)

C*-94/P.3514 CHODA M. Extension Algebras of II_{1} Factors via *-Endomorphisms Dpt. Math., Osaka Kyoiku Univ., Asahigaoka, Kashiwara 582 (Japan)

C*-94/P.3515 CHODA M. Square Roots of the Canonical Shifts Dpt. Math., Osaka Kyoiku Univ, Tennoji, Osaka 543 (Japan)

C*-94/P.3516 CHODA M. Canonical *-Endomorphisms and Simple C*-Algebras Dpt. Math., Osaka Kyoiku Univ., Asahigaoka, Kashiwara 582 (Japan)

C*-94/P.3517 CHODA M.

Conjugate but Non Inner Conjugate Subfactors Dpt. Math., Osaka Kyoiku Univ., Asahigaoka, Kashiwara 582 (Japan)

C*-94/P.3518 BEKKA M.E.B., VALETTE A. Group Cohomology, Harmonic Functions and the First L^{2}-Betti Number Dpt. Math., Univ. Metz, Ile du Saulcy, F-57045 Metz (France)

C*-94/P.3519 RORDAM M. A Short Proof of Elliott’s Theorem : O_{2} ⨂ O_{2} ≅ O_{2} Inst. Mat. of Dat., Odense Univ., Campusvej 55, DK-5230 Odense M (Denmark) E-MAIL : mikael@imada.ou.dk

C*-94/P.3520 BAAJ S. Representation Reguliere du groupe quantique des deplacements de Woronowicz Dpt. Math. Inform., Univ. d’Orleans, 45046 Orleans (France)

C*-94/P.3521 VAN DAELE A. The Haar Measure on a Compact Quantum Group Dpt. Math., K.U. Leuven, Celestijnenlaan 200B, B-3001 Heverlee (Belgium)

C*-94/P.3522 KAWAHIGASHU Y. Orbifold Subfactors, Central Sequences and the Relative Jones Invariant khi Dpt. Math. Sciences, Univ. Tokyo, Hongo, Tokyo 113 (Japan) E-MAIL : yasuyuki@tansei.cc.u-tokyo.ac.jp

C*-94/P.3523 KUSTERMANS J. Two Notions of Discrete Quantum Groups Dpt. Math., K.U. Leuven, Celestijnenlaan 200B, B-3001 Heverlee (Belgium)

C*-94/P.3524 VAN DAELE A., WORONOWICZ S.L. Duality for the Quantum E(2) Group Dpt. Math., K.U. Leuven, Celestijnenlaan 200B, B-3001 Heverlee (Belgium)

C*-94/P.3525 VAN DAELE A. Discrete Quantum Groups Dpt. Math., K.U. Leuven, Celestijnenlaan 200B, B-3001 Heverlee (Belgium)

C*-94/P.3526 POPA S. Classification of Subfactors of Finite Depth of the Hyperfinite Type III_{1} Factor Math. Dept., University of Calif. of Los Angeles (UCLA), Los Angeles CA90034-1555 (USA)

C*-94/P.3527 POPA S. Symmetric Enveloping Algebras, Amenability and AFD Properties for Subfactors Math. Dept., University of Calif. of Los Angeles (UCLA), Los Angeles CA90034-1555 (USA)

C*-94/P.3528 RORDAM M. Classification of Certain Infinite Simple C*-Algebras, III Inst. Mat. of Dat., Odense Univ., Campusvej 55, DK-5230 Odense M (Denmark) E-MAIL : mikael@imada.ou.dk

C*-94/P.3529 NIELSEN K.E., THOMSEN K. Limits of Circle Algebras Aarhus Univers., Mat. Inst., Ny Munkegade, 8000 Aarhus C (Denmark) E-MAIL : matkt@mi.aau.dk

C*-94/P.3530 MORITA A., EILERS S., EXEL R. Morita Equivalence for Crossed Products by Hilbert C*-Bimodules Matematisk Inst., Kopenhavns Universitet, Universitetsparken 5, 2100 Copenhagen 0 (Denmark) E-MAIL : abadi@cmat.edu.uy

eilers@math.ku.dk exel@ime.usp.br

C*-94/P.3531 HAAGERUP U., STORMER E. Positive Projections of von Neumann Algebras onto JW-Algebras Inst. Mat. of Dat., Odense Univ., Campusvej 55, DK-5230 Odense M (Denmark) E-MAIL : haagerup@imada.ou.dk

C*-94/P.3532 ANGELESCU N., VERBEURE A. Variational Solution of a Superfluidity Model Inst. voor Theor. Fysica, K.U. Leuven, Celestijnenlaan 200 D, B-3001 Leuven (Belgium)

C*-94/P.3533 ANGELESCU N., BUNDARU M., PULE J.V. Large Deviations in a Semi-Infinite Layered Ferromagnet with Mean-Field Interactions :Classical and Quantum Models Inst. voor Theor. Fysica, K.U. Leuven, Celestijnenlaan 200 D, B-3001 Leuven (Belgium)

C*-94/P.3534 WU LIANG SEN Split Inclusion and Metrically Nuclear Map Dpt Math., East China Normal Univ., Shanghai 200062, (P.R. China)

C*-94/P.3535 WU LIANG SEN Split Inclusion and Hilbert Space Factorable Maps Dpt Math., East China Normal Univ., Shanghai 200062, (P.R. China)

C*-94/P.3536 JI RONGHUI, SCHWEITZER L.B. Spectral Invariance of Smooth Crossed Products, and Rapid Decay Locally Compact Groups Dpt Math. Sciences, IUPUI, 402 N. Blackford Street, Indianapolis, IN 46202-3216(USA) E-MAIL : ronji@math.iupui.edu

schweitzer@math.la.asu.edu

C*-94/P.3537 CHU CHO-HO, LAU KA-SING Solutions of the Operator-Valued Integrated Cauchy Functional Equation Goldsmiths’ College, Univ. London, London SE14 (GB)

C*-94/P.3538 KOSAKI K., MUNEMASA A., YAMAGAMI S.

On Fusion Algebras Associated to Finite Group Actions Graduate School of Math., Kyushu Univ., Ropponmatsu, Fukuoka 810 (Japan)

C*-94/P.3539 LAPIDUS M.L., PANG M.M.H. Eigenfunctions of the Koch Snowflake Domain Dpt. Math., Sproul Hall, The University of California, Riverside CA 92521-0135 (USA) E-MAIL : lapidus@math.ucr.edu

pang@math.cornell.edu

C*-94/P.3540 LAPIDUS M.L. Analysis on Fractals, Laplacians on Self-Similar Sets, Noncommutative Geometry and Spectral Dimensions Dpt. Math., Sproul Hall, The University of California, Riverside CA 92521-0135 (USA) E-MAIL : lapidus@math.ucr.edu

C*-94/P.3541 LAPIDUS M.L., POMERANCE C. Counterexamples to the Modified Weyl-Berry Conjecture on Fractal Drums Dpt. Math., Sproul Hall, The University of California, Riverside CA 92521-0135 (USA) E-MAIL : lapidus@math.ucr.edu

carl@ada.math.uga.edu

C*-94/P.3542 CUNTZ J., QUILLEN D. Excision in Bivariant Periodic Cyclic Cohomology Math. Inst., Univ. Heidelberg, Im Neuenheimer Feld 288, 69120 Heidelberg (RFA)

C*-94/P.3543 FREDENHAGEN K., GABERDIEL M.R.,

RUGER S.M.

Scattering States of Plektons (Particles with Braid Group Statistics) in 2+1 Dimensional Quantum Field Theory. II Inst. fur Theor. Physik, Univ. Hamburg, D 22761 Hamburg (RFA)